OpenStudy (anonymous):

Solve the equation using factoring. x2 + 8x - 20 = 0 A. {-2, -10} B. {2, 10} C. {-2, 10} D. {2, -10}

OpenStudy (mysesshou):

(x-2)(x+10) (x-2) = 0 and (x+10)=0 x=2, x=-10

OpenStudy (mysesshou):

x2 + 8x - 20 = 0 (-2)^2 + 8(-2) - 20 = 4 -16-20 = -32 (-10)^2 + 8(-10) - 20 = 100 - 80 - 20 = 0 (2)^2 + 8(2) - 20 = 4 + 16 - 20 = 0 I still get D. Someone pls message me to explain further. I'm confused why it is A.

OpenStudy (across):

\[x^2+8x-20=0\implies(x+10)(x-2)=0\implies x=-10\text{ or }2.\]mysesshou is right.

OpenStudy (mysesshou):

AlexisDelgado, we support D, with proven examples. The End. Bye everyone :)

OpenStudy (anonymous):

They're not. Sorry about that

OpenStudy (amistre64):

x2 + 8x - 20 = 0 that -20 tells us we have different signed terms +8 means that the smaller root is positive

OpenStudy (anonymous):

It is A. Here is my work.

OpenStudy (amistre64):

lol, it does look like something id work out :)

OpenStudy (amistre64):

check your sign in step 35